OpenStudy (anonymous):

is the sequence arithmetic? if so, identify the common difference. 17, 22, 33, 46

OpenStudy (helder_edwin):

for the squence to be arithmetic the differerence of any term and the term before it has to be constant

OpenStudy (anonymous):

alright so is it?

OpenStudy (helder_edwin):

u have a_1=17 a_2=22 and a_3=33 but a_2-a_1=22-17=5 a_3-a_2=33-22=11

OpenStudy (anonymous):

so its not consitant

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

its not. right?

OpenStudy (helder_edwin):

yes. it is not an arithmetic sequence.

OpenStudy (anonymous):

alright :-) thank you so much

OpenStudy (helder_edwin):

u r welcome